OpenStudy (sambot):

28x12365 help? i will fan

OpenStudy (anonymous):

346220

OpenStudy (sambot):

thanks

OpenStudy (anonymous):

medal?

OpenStudy (sambot):

there

rebeccaxhawaii (rebeccaxhawaii):

are u jokin

OpenStudy (sambot):

no...sorry

OpenStudy (anonymous):

346220

OpenStudy (crushcandy):

\[2*12365=346220\]

OpenStudy (anonymous):

28*12365=(30-2)*12365=30*12365-2*12365 =370950-24730 =346220